OpenStudy (anonymous):

What is the equation of this circle in standard form? (x − 5.5)2 + (y − 4)2 = 7 (x + 5.5)2 + (y + 4)2 = 3.5 (x − 5.5)2 + (y − 4)2 = 3.5 (x + 5.5)2 + (y + 4)2 = 12.25 (x − 5.5)2 + (y − 4)2 = 12.25

OpenStudy (anonymous):



OpenStudy (jdoe0001):

use distance formula to find the radius and the midpoint formula to find the center then plug them in the standard equation

OpenStudy (anonymous):

Is the answer E?

OpenStudy (jdoe0001):

\(\bf (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2=r^2 \\ \quad \\ \quad \\ \textit{distance between 2 points}\to radius\to r\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ M&({\color{red}{ 2}}\quad ,&{\color{blue}{ 4}})\quad N&({\color{red}{ 9}}\quad ,&{\color{blue}{ 4}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2} \\ \quad \\ \textit{middle point of 2 points }\to center\to ({\color{brown}{ h,k}})\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ M&({\color{red}{ 2}}\quad ,&{\color{blue}{ 4}})\quad N&({\color{red}{ 9}}\quad ,&{\color{blue}{ 4}}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\)

OpenStudy (jdoe0001):

\(\bf (x-{\color{brown}{ h}})^2+(y-{\color{brown}{ k}})^2=r^2 \\ \quad \\ \quad \\ \textit{distance between 2 points}\to radius\to r\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ M&({\color{red}{ 2}}\quad ,&{\color{blue}{ 4}})\quad N&({\color{red}{ 9}}\quad ,&{\color{blue}{ 4}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2} \\ \quad \\ \textit{middle point of 2 points }\to center\to ({\color{brown}{ h,k}})\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ M&({\color{red}{ 2}}\quad ,&{\color{blue}{ 4}})\quad N&({\color{red}{ 9}}\quad ,&{\color{blue}{ 4}}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\) anyhow, there

OpenStudy (anonymous):

Thank you

OpenStudy (jdoe0001):

yw